OpenStudy (anonymous):

For N0 x e^(kt). If the constant k = (ln20/18) and the initial No of cases was estimated to be 50,000 in the year 1987, how many cases will there be in the year 2000, in terms of ln?

OpenStudy (anonymous):

1987, t=0.

OpenStudy (anonymous):

then, get the difference 2000 - 1987 = _______?

OpenStudy (anonymous):

that will be the value of t, then plug-in the values now. that's it:)

OpenStudy (anonymous):

okay, thanks. I came up with 50000 e^(ln20/18) (13) However, after the e/ln cancels out..I'm not sure if I should multiply 50000 x 20 x 13/18 ? or is it 50000 x 20 6 (13/18)?

OpenStudy (anonymous):

@UnkleRhaukus @violy

OpenStudy (anonymous):

note you will have, 50000e^ln((20/18)(13)).

OpenStudy (anonymous):

just multiply across. remember to use PEMDAS as well.

OpenStudy (anonymous):

Ah, I see! Thanks!!! @violy

OpenStudy (unklerhaukus):

\[50000 e^{\ln\frac{20}{18} (13)}=50000 e^{13\ln\frac{20}{18} }=50000 e^{{\ln\left(\frac{20}{18}\right)^{13}}}\]

OpenStudy (anonymous):

thanks @UnkleRhaukus

OpenStudy (anonymous):

So if I were to simplify the equation below. and solve for t.\[ 100=e ^ (\ln20/18)t \] how would I solve for t in terms of ln? if e ^(ln20/18)(t) @UnkleRhaukus?? thanks!

OpenStudy (anonymous):

would I cross multiply again?

OpenStudy (unklerhaukus):

\[100= e ^{\ln\frac{20}{18}t}=\left(\frac{20}{18}\right)^t\]

OpenStudy (anonymous):

Ah, I see. And if I want to solve for t after that, how would I do that??

OpenStudy (unklerhaukus):

um

OpenStudy (unklerhaukus):

\[100=\left(\frac{5}{4}\right)^t\] \[\ln_{5/4}100=t\]

OpenStudy (anonymous):

Oh wow!! Great!!! thanks so much!!

OpenStudy (unklerhaukus):

ops made a mistakes

OpenStudy (anonymous):

what mistake?

OpenStudy (unklerhaukus):

20/18≠5/4

OpenStudy (anonymous):

Oh ok! ...so it will be ln(20/18)100

OpenStudy (unklerhaukus):

can you simplify that ?

OpenStudy (unklerhaukus):

\[N(t)=N_0 \times e^{kt}\]\[\frac{N(t)}{N_0}= e^{kt}\]\[\ln\frac{N(t)}{N_0}=kt\]\[\frac{\ln\frac{N(t)}{N_0}}k=t\]

OpenStudy (anonymous):

Ok thanks!! A lot to chew on, but it is making a lot more sense now!! :)

OpenStudy (unklerhaukus):

this is first order growth equation, occurs all kinds of places, if k was negative it would the first order decay which also occurs lots of places, so it is good to lean this equation because you will encounter it time and time again

OpenStudy (anonymous):

thanks!! I'll definitely keep working on it!